# Double Privacy Guard: Robust Traceable Adversarial Watermarking against Face Recognition

### v1

@@ V-D Visualization and Analysis
DPG 0.326 0.903 0.810 0.896 0.381 0.911 0.716 0.911 0.643 0.898 0.355 0.913
TABLE V: Ablation Experiment. w/o Meta indicates DPG is trained with a direct ensemble of FR models, without meta-optimization.
Figure 5: Visualization using Grad-CAM produces attention maps on CosFace model and IRSE50 model.
V-E Ablation Study
To illustrate the influence of incorporating the meta-optimization strategy compatible with multiple tasks in end-to-end optimization on the comprehensive performance of DPG, we present the results of ablation experiments in Table V. Meta-optimization strategy in DPG simultaneously enhances image quality and ASR. This is attributed to the encoder quickly adapting to new tasks with minimal gradient updates during meta-optimization, acquiring universal adversarial features for the carrier against unknown victim models. This avoids excessive reliance on penalizing image quality to enhance adversarial characteristics.
VI Conclusion
To address the limitations of existing privacy protection methods, we propose DPG, the first method that protects facial image privacy from both adversarial and traceable perspectives. We first propose an information-guided adversarial attack strategy, wich guides recognizable features of the image to deviate from the source identity features through an identity-specific watermark. Then we introduce a meta-optimization strategy that is compatible with multiple tasks to solve the optimization conflict between transferability and watermark invisibility. Experiments confirm the significant transferability, traceability and robustness of DPG on the SOTA FR models.

### v2

@@ 5.4 Visualization and Analysis
DIP-Watermark 0.326 0.903 0.810 0.896 0.381 0.911 0.716 0.911 0.643 0.898 0.355 0.913
TABLE V: The effectiveness evaluation of meta-optimization methods. w/o Meta indicates DIP-Watermark is trained with a direct ensemble of FR models, without meta-optimization.
  $\lambda_{adv}:\lambda_{en}:\lambda_{wm}$ IR152 IRSE50 IResNet100-Arc MobileFace FaceNet IResNet100-Cos
ASR ACC ASR ACC ASR ACC ASR ACC ASR ACC ASR ACC
1:1:1 0.025 0.932 0.125 0.954 0.005 0.987 0.006 0.977 0.000 0.983 0.000 0.967
10:1:1 0.100 0.977 0.092 0.968 0.000 0.972 0.167 0.983 0.063 0.960 0.000 0.968
100:0.05:0.05 0.326 0.992 0.810 0.997 0.381 0.996 0.716 0.997 0.643 0.994 0.355 0.996
TABLE VI: Impact of various weight coefficients on model performance.
Figure 5: Visualization using Grad-CAM produces attention maps on IResNet100-Cos model and IRSE50 model.
5.5 Ablation Study
To illustrate the influence of incorporating the meta-optimization strategy compatible with multiple tasks in end-to-end optimization on the comprehensive performance of DIP-Watermark, we present the results of ablation experiments in Table V. Meta-optimization strategy in DIP-Watermark simultaneously enhances image quality and ASR. This is attributed to the encoder quickly adapting to new tasks with minimal gradient updates during meta-optimization, acquiring universal adversarial features for the carrier against unknown victim models. This avoids excessive reliance on penalizing image quality to enhance adversarial characteristics.
Additionally, we evaluate the impact of different weights on model performance, as shown in Table VI. As previously discussed, under equal weights, adversarial loss fails to converge after uniform training, resulting in poor black-box adversarial performance across all models. However, with the introduction of magnitude disparities, the training loss effectively converges, significantly enhancing adversarial performance, indicating the necessity of uniform loss value magnitudes during initial training stages.
6 Conclusion
To address the limitations of existing privacy protection methods, we propose DIP-Watermark, the first method that protects facial image privacy from both adversarial and traceable perspectives. DIP-Watermark embeds a robust watermark into facial images, thwarting unauthorized malicious FR retrievals while enabling authorizers to extract the watermark for original identity tracing, thus mitigating potential threats of privacy leakage and identity impersonation. Specifically, we first propose an information-guided adversarial attack strategy, wich guides recognizable features of the image to deviate from the source identity features through an identity-specific watermark. Then we introduce a meta-optimization strategy that is compatible with multiple tasks to solve the optimization conflict between transferability and watermark invisibility. We conducted extensive experiments on two large-scale facial datasets, confirming that DIP-Watermark achieves significant adversarial transferability, traceability, and robustness against state-of-the-art face recognition models. It surpasses both adversarial attack-based methods and methods that simply combine adversarial attacks directly with watermarking.
Acknowledgments
This work was supported by National Natural Science Foundation of China NSFC (No. 62072343); the Fundamental Research Funds for the Central Universities (No. 2042023kf0228); the National Key Research and Development Program of China (No. 2019QY(Y)0206).